### v1

# SwapAnything: Enabling Arbitrary Object Swapping in Personalized Visual Editing

Abstract
Effective editing of personal content holds a pivotal role in enabling individuals to express their creativity, weaving captivating narratives within their visual stories, and elevate the overall quality and impact of their visual content. Therefore, in this work, we introduce SwapAnything, a novel framework that can swap any objects in an image with personalized concepts given by the reference, while keeping the context unchanged. Compared with existing methods for personalized subject swapping, SwapAnything has three unique advantages: (1) precise control of arbitrary objects and parts rather than the main subject, (2) more faithful preservation of context pixels, (3) better adaptation of the personalized concept to the image. First, we propose targeted variable swapping to apply region control over latent feature maps and swap masked variables for faithful context preservation and initial semantic concept swapping. Then, we introduce appearance adaptation, to seamlessly adapt the semantic concept into the original image in terms of target location, shape, style, and content during the image generation process. Extensive results on both human and automatic evaluation demonstrate significant improvements of our approach over baseline methods on personalized swapping. Furthermore, SwapAnything shows its precise and faithful swapping abilities across single object, multiple objects, partial object, and cross-domain swapping tasks. SwapAnything also achieves great performance on text-based swapping and tasks beyond swapping such as object insertion.
Figure 1: SwapAnything results on various personalized image swapping tasks. SwapAnything is adept at precise, arbitrary object replacement in a source image with a personalized reference, and achieves high-fidelity swapping results without influencing any context pixels. We demonstrate its general effect in single-object, multi-object, partial-object, and cross-domain swapping tasks.
1 Introduction
In today’s digital age marked by the prolific creation and widespread sharing of personal content, generative models (Rombach et al., 2022; Crowson et al., 2022; Ding et al., 2021; Blattmann et al., 2022) have risen as a potent tool for self-expression, storytelling, and amplifying the impact of visual narratives. Among existing image editing methods (Meng et al., 2022; Hertz et al., 2022; Chen et al., 2023b) for content creation that empowered individuals to convey their creative instincts, weave captivating narratives into their visual stories, and enhance the quality of their visual representations, personalize content swapping (Gu et al., 2023a; Li et al., 2023b; Gu et al., 2023b), which targets at creating and compositing new images with user-specified visual concept, attracted significant interest due to its wide-ranging applications in E-commerce, entertainment, and professional editing.
Achieving arbitrary personalized content swapping necessitates a deep understanding of the visual concept inherent to both the original and replacement subjects. There are several crucial challenges. Firstly, arbitrary swapping demands significantly greater flexibility from the swapping technique compared with swapping the main subject, due to the varied nature of the content being exchanged. Secondly, an ideal swap requires flawless preservation of the surrounding context pixels, ensuring that only the designated target area undergoes modification. The third challenge lies in accurately integrating the personalized content into the target image in a harmonious manner while preserving the original poses and styles.
Existing works often fall short of addressing these challenges. Most of existing research (Ruiz et al., 2023; Shi et al., 2023; Chen et al., 2023b; Jia et al., 2023; Wang et al., 2024) are focused on personalized image synthesis, which seeks to create new images with personalized content. Although these approaches can synthesize high-fidelity content, they cannot edit or replace the visual content in an existing image.  Li et al. (2023b); Cao et al. (2023); Meng et al. (2022); Avrahami et al. (2023) have tried to remove and regenerate the object via masks, they often struggle to adapt the new concept into the image. Recently studies (Patashnik et al., 2023; Gu et al., 2023a; Tumanyan et al., 2023) focus on object swapping and replacement by tweaking intermediate variables affecting the object’s features. However, this approach lacks the precision needed for localized object swapping, resulting in limited visual qualities. Besides, those methods mainly work on main subject swapping, and they are incapable of arbitrary object swapping.
To address these challenges, we introduce SwapAnything, a framework that utilizes pre-trained diffusion models to streamline personalized arbitrary object swapping. Unlike previous work, our work is designed for arbitrary swapping tasks with perfect context pixel preservation and harmonious object transition. Our method begins by exploring an informative representation of the source image on a diffusion model. We found that various variables in the diffusion process especially latent features from U-net have a correspondent relation with the image. So we propose to keep the context pixels in the source image by preserving the correspondent part in those variables in the swapping process. This process is tailored to precisely swap specific areas, ensuring the preservation of other objects and the background’s integrity. The object information in the source image is also selected for appearance adaptation. More specifically, location adaptation controls the location where the new concept should be swapped. Style adaptation ensures stylistic harmony between the concept object and the original image, fostering a natural and cohesive visual presentation. Additionally, scale adaptation is introduced to modulate the target object’s shape, ensuring its congruence with the spatial and dimensional aspects of the source image. Last, content adaptation is crucial for smoothly generating the new concept, enabling a seamless blend that mitigates any artifacts or unnatural transitions. With these specialized adaptations, SwapAnything provides a heightened level of precision and refinement in the realm of object-driven image content swapping as shown in Fig. 1.
Our main contributions are: 1) We propose SwapAnything, a general framework for both personalized swapping and text-based swapping on single object, multiple objects, partial object, and cross-domain object. 2) We identified key variables for content preservation and proposed targeted swapping for perfect background preservation. 3) We designed a sophisticated appearance adaptation process to adapt the concept image into the source object. 4) Our approach has demonstrated exceptional performance through comprehensive qualitative assessments and quantitative analyses on swapping tasks and tasks beyond swapping such as insertion.
2 Related Work
2.1 Text-guided Image Editing
With recent progress in diffusion model, text-guided Image editing has been largely explored (Avrahami et al., 2023). Image editing driven by text aims to alter an existing image following the textual guidelines provided, while ensuring certain elements or features of the original image remain unchanged. Initial efforts using GAN models (Karras et al., 2019) were restricted to specific object domains. However, diffusion-oriented techniques (Zhang et al., 2023; Nichol et al., 2022; Feng et al., 2023) have surpassed this limitation, enabling text-driven image modifications. While these techniques produce captivating outcomes, many grapple with localized edits. As a result, manual masks (Meng et al., 2022; Zeng et al., 2023; Meng et al., 2022) are often help to delineate the editing zones. Though employing cross-attention (Hertz et al., 2022) or spatial attributes (Tumanyan et al., 2023) can achieve localized edits, they face challenges with flexible changes (like altering poses) and preserving the initial image composition.
2.2 Object Driven Image Editing
In object-driven image generation, the object could be inverted into the texual space so to represented by a new token such as “***”  (Gal et al., 2023; Ruiz et al., 2023; Chen et al., 2023a; Shi et al., 2023; Tewel et al., 2023). Image editing guided by exemplars spans a vast array of uses. Much of the research in this area (Wang et al., 2019; Huang et al., 2018; Zhou et al., 2021) falls under the umbrella of example-based image transformation tasks. These tasks leverage various sources of information, from stylized images (Liu et al., 2021; Deng et al., 2022; Zhang et al., 2022) and layouts (Yang et al., 2023b; Li et al., 2023c; Jahn et al., 2021), to skeletons (Li et al., 2023c) and sketches or outlines (Seo et al., 2023). Given the versatility of stylized images, the art of image style adaptation (Liao et al., 2017; Zhang et al., 2020) has garnered significant interest. These methods primarily focus on establishing a detailed match between the input and reference visuals but falter when it comes to localized alterations. To facilitate localized changes, especially with flexible transformations, tools like bounding boxes and skeletons have been introduced. However, these tools often demand manual input, which can be challenging for users. A novel approach (Yang et al., 2023a) frames exemplar-driven image editing as a task of filling in gaps, maintaining the context while transferring semantic elements from the reference to the original image. While DreamEdit (Li et al., 2023b) employs a step-by-step gap-filling method for object substitution, it doesn’t effectively bridge the connection between the original and desired objects. In contrast, our method ensures that crucial features, like body movements and facial expressions, stay consistent. Meanwhile, Photoswap (Gu et al., 2023a), CustomEdit (Choi et al., 2023) and BLIP-Diffusion (Li et al., 2023a) achieved personalized object swapping. However, it does not keep the non-object pixel background intact. While our method directly focuses on the editing area without influencing other objects and background.
3 Preliminary
Diffusion Models belong to the family of generative models that are based on stochastic processes. They generate an image by iteratively reducing noise from an initial distribution. Starting from a point of random noise denoted as ${\boldsymbol{z}}_{T}$, which follows a normal distribution ${\boldsymbol{z}}_{T}\sim\mathcal{N}(0,\mathbf{I})$, the diffusion model denoises each instance ${\boldsymbol{z}}_{t}$, thus producing ${\boldsymbol{z}}_{t-1}$. Diffusion models predict and reverse the noise at each step in the diffusion sequence to arrive at the final denoised image.
In our research, we employ the pre-trained text-to-image diffusion framework Stable Diffusion Rombach et al. (2022). This model encodes images into a latent space and incrementally denoises the encoded latent representation. The Stable Diffusion model operates on a U-Net architecture (Ronneberger et al., 2015), where the latent representation ${\boldsymbol{z}}_{t-1}$ at any given step is derived from the text prompt $P$ and the previous latent state ${\boldsymbol{z}}_{t}$, as indicated by the following equation:
${\boldsymbol{z}}_{t-1}={\boldsymbol{\epsilon\theta}}({\boldsymbol{z}}_{t},P,t)$ (1)
This U-Net includes sequence of layers that repeatedly apply self-attention and cross-attention mechanisms. In self-attention, the latent image feature $z_{t}$, is first projected into query $Q_{self}$, $K_{self}$, $V_{self}$, which will be used to get self-attention map ${\boldsymbol{A}}$ and self-attention output ${\boldsymbol{\phi}}$.
$\begin{split}M&=\text{softmax}\left(\frac{Q_{self}\cdot K_{self}^{T}}{\sqrt{d}}\right)\\ \phi&=M\cdot V_{self}\end{split}$ (2)
Meanwhile, for cross-attention layer, the feature out of previous self-attention layer is projected into $Q_{cross}$, while feature embedding of textual prompt is projected into $K_{cross}$ and $V_{cross}$.
$\begin{split}A&=\text{softmax}\left(\frac{Q_{cross}\cdot K_{cross}^{T}}{\sqrt{d}}\right)\end{split}$ (3)
where $A$ is the cross-attention map. In this work, we study the swapping of $A$, $M$, $\phi$ and $z$.
Figure 2: Overview of SwapAnything on swapping a object from a source image ($I_{src}$) into a personalized concept (<*><{*}>< * >) to get the target image ($I_{target}$). The personalized concept is first converted into textual space to be treated as concept appearance. Meanwhile, the source image is first inverted into initial noise to obtain U-Net variables (including latent feature, attention map, and attention output). Targeted variable swapping preserves the context pixels in the source image. The appearance adaptation process then utilizes these informative variables to integrate the concept into the target image.
4 SwapAnything
Figure 3: Swapping process in SwapAnything. The left part shows the correspondence between latent feature $z$ and the Generated image. The right part shows the procedure of targeted variable manipulation in the U-Net diffusion process.
In this section, we introduce the SwapAnything framework that uses a diffusion model to swap on targeting area faithfully while keeping the context pixels unchanged. Fig. 2 illustrates our framework’s overall structure. For source image $I_{src}$, we first invert it to a latent noise and then obtain the feature representations $V_{src}$, which will be used during the target image $I_{src}$ generation process. In Sec. 4.1, we discuss how to preserve the non-target pixels in the source image perfectly, and how to select and transfer key information about the source image. Following this, in Sec. 4.2, we introduce the appearance adaptation pipeline that uses the key information to integrate the new concept into the source image seamlessly.
4.1 Targeted Variable Swapping
Intermediate variables in the U-Net of a diffusion model have been proven informative about the content of the generated image (Gu et al., 2023a; Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023). They usually focus on the study of variables inside of U-net structure such as attention map, and attention output, while the output of U-Net at each diffusion step, i.e., latent image feature ${\boldsymbol{z}}$ is not widely explored before. We argue that the latent image feature ${\boldsymbol{z}}$ contains more information on image content control. The image generation process for the latent diffusion model is achieved by denoising the ${\boldsymbol{z}}$ to arrive at a clear representation of a high-quality image, whereas all other variables inside of U-net indirectly affect the image by impacting ${\boldsymbol{z}}$. In contrast to simply swapping ${\boldsymbol{z}}$ like other variables, which would erase the new image’s unique details and result in a mere duplication of the original image, our investigation revealed a significant correlation between the latent feature ${\boldsymbol{z}}$ and the produced image, including a pixel-level correspondence. In the left part of Fig. 3, we visualize the main component of the averaged latent feature $z$ across all diffusion steps. The finding that It has a part-to-part correspondence with the generated images indicates the potential of localized editing by manipulating the latent feature.
Consequently, we consider a strategy where only the context pixels within ${\boldsymbol{z}}$ are altered, affecting solely the intended pixel. Although this method yields suboptimal outcomes, akin to merging two unrelated images, it prompts us to devise two remedial measures. Firstly, we suggest limiting the exchange of the latent feature to the initial stages of diffusion, allowing subsequent steps to smooth out any discordance in the latent space. Furthermore, our exploration into Unet’s cross-attention map ${\boldsymbol{M}}$, self-attention map ${\boldsymbol{A}}$, and self-attention output ${\boldsymbol{\psi}}$ reveal their potential in mitigating artifacts. Swapping those facilitates the alignment of the latent features between the source and target images before the partial-swapping between them. In short, all the variables mentioned above in both the source image and target image generation process could be resized into the shape of the mask, where the mask can be utilized for the swapping process.
$V_{target}=g(f(V_{src})*(1-\mathbf{M_{src}}{})+f(V_{target})*\mathbf{M_{src}}{})$ (4)
Here $V$ includes latent feature $z$, and other assistant variable cross-attention map ${\boldsymbol{M}}$, self-attention map ${\boldsymbol{A}}$, and self-attention output ${\boldsymbol{\phi}}$. $f(\cdot)$ means the transformation process to the shape of the mask, while $g(\cdot)$ means the transformation back to the original space. For simplicity, we ignore all $f(\cdot)$ and $g(\cdot)$ in the following text. The content in the latent feature of the source image is changing as the diffusion process continues. Therefore, the location of the correspondent pixel in latent space may change over diffusion steps. A direct solution is to decode the latent feature ${\boldsymbol{z}}$ into an image at each step and extract the mask dynamically according to the object location in the generated image. However, we find that a changing mask usually confuses the model and leads to a less optimal performance. Therefore, while using the same high-quality mask through the diffusion process. We find that the mask could either be extracted from the source image directly using an off-the-shelf model or from the generation process as in Simsar et al. (2023); Patashnik et al. (2023). Please check the Appendix for more details.
4.2 Appearance Adaptation
In this section, we introduce the appearance adaptation process that adapts the concept into the source image, which necessitates meticulous adjustments across several dimensions: location, style, scale, and content. Our framework enhances realism and coherence in image manipulation, marking a significant advancement in the field.
4.2.1 Location Adaptation
Various intermediate variables have been proven to correlate with the final generated image. Although with great performance, it did not achieve local swap and thus the background is modified inevitably. As shown in Fig. 2, for each step, instead of directly swapping the whole variable, we conduct local swapping to only swapping the non-object position. Also, to enhance the swapping results, we further propose to conduct local swapping on the latent representation ${\boldsymbol{z}}$ directly. $\mathbf{M_{src}}{}$ is a 2-dimension variable containing 0 and 1. It is the same size as of the source image and value 1 marks the swapping location. To simplify the expression, here we denote three U-Net variables attention map, attention output, and latent representation for the original image recovery process as $V_{src}$, denote the ones generated via target text prompt as $V_{concept}$, then we define the target variable used as $V_{target}^{bg}$ as the background information of the target variable, which can be obtained from Eq. 5.
$V_{target}^{bg}=V_{src}*(1-\mathbf{M_{src}}{})$ (5)
The non-masked area is the swapping target area, where the variable will be generated via the target text prompt to inject the concept appearance. Location adaptation extends beyond the object swapping tasks; we also discovered its profound capability for object insertion. For further details and results, please refer to the Appendix.
4.2.2 Style Adaptation
An ideal object swapping should keep the style unchanged. The object information in the generated variables is injected via the new concept token. Some style attributes could be already bound with the token. Therefore, solely generating the foreground information via the text prompt might lead to style inconsistency. Recently, Karras et al. (2019); Park et al. (2019) found that adding such normalization layers can help improve the conditional image generation quality because such activation functions modulation. Unlike from them, in our work, we employ the AdaIN (Adaptive Instance Normalization) to modulate the swapping features with spatial constraints. We follow Eq. 6 and Eq. 7 to denormalize the $V_{concept}$ with the mean and variance from $V_{src}$ in each time step for $V_{target}$ during the image generation process. As a result, we find that by modulating the concept feature, the generated content can adaptively follow the original style in the source image.
$V_{concept}^{{}^{\prime}}=MaskedAdaIN(V_{scr},V_{concept},M_{src})$ (6)
$V_{target}^{fg}=V_{concept}^{{}^{\prime}}*M_{src}$ (7)
MaskedAdaIN utilizes the mean and variance from the masked region in the AdaIN calculation. Then we have the blended feature representations for $V_{target}$:
$V_{target}=V_{target}^{fg}+V_{target}^{bg}$ (8)
4.2.3 Scale Adaptation
The proportion of an object compared to its environment and other elements in the image is crucial for coherence. A swapping result with improper scaling can disturb the aesthetic balance, resulting in a disjoint appearance of the image.
Guidance from an external classifier in the inference process of Diffusion models could influence the diffusion noise to control the generated image. Epstein et al. (2023) shows that the guidance can also be used on the attention map to control the generation. Similarly, we adapt the mask guidance Eq. 9 to better align the shape between the source object and the target object.
$\hat{\epsilon}_{t}$ $=(1+s)\epsilon_{\theta}(z_{t};t,y)-s\epsilon_{\theta}(z_{t};t,\emptyset)$
$\quad+v\sigma_{t}\nabla_{z_{t}}\|M_{src}-Shape(M_{src})\left(k\right)\|_{1}$ (9)
where $s$ is the classifier-free guidance strength and $v$ is an additional guidance weight for $g$. As with classifier guidance, we scale by $\sigma_{t}$ to convert the score function to a prediction of $\mathbf{\epsilon}_{t}$. $Shape(\mathbf{M_{src}}{})$ denotes the object shape as identified in the attention layer. Here the energy function $g$ is set as $\|\mathbf{M_{src}}{}-Shape(\mathbf{M_{src}}{})\|_{1}$ to calculate the shape difference between the original object mask and the extracted shape in the attention layer, which indicates the deviation between the ideal shape and shape during the diffusion process.
4.2.4 Content Adaptation
A binary mask without smoothing has a high-frequency transition at the edge — it jumps abruptly from 0 to 1. When used to merge two intermediate variables from two different diffusion processes, this can result in high-frequency artifacts at the boundary, such as jagged edges or a halo effect. Smoothing the mask transitions these high frequencies into lower frequencies, which blends the images more naturally and eliminates such artifacts. A smooth mask creates a feathering effect at the edges of the transition. This makes the merged area appear more coherent as if the two images naturally blend into each other rather than being cut off abruptly. Therefore, for the diffusion process, we specifically design two masks according to the feature of diffusion models.
Without this smoothing, the boundary between the images would be sharply defined, leading to a jarring and unnatural appearance. The Gaussian Blur softens the edges, blending the images more seamlessly. To augment this improvement, we introduce two smoothing techniques for binary masks, applied across both spatial dimensions and temporal steps. These techniques serve to refine the swapping process, mitigating artifacts and ensuring a smoother, more natural integration of the swapped regions. This results in an enriched visual output, seamlessly blending the inserted objects or object parts into the overall image composition.
Linear Boundary Interpolation: This is a process where the sharp transition between the area with 1s and the area with 0s in your binary array is made gradual. One way to achieve this is by using a convolution with a smoothing kernel (like a Gaussian kernel) that will average the values in the vicinity of each point, effectively creating a gradient at the boundary.
$\delta(\mathbf{M_{src}}{})$ $=\mathbf{M_{src}}{}\oplus K$
$S$ $=\delta(\mathbf{M_{src}}{})\ast G$
$S^{\prime}[i,j]$ =$=\begin{cases}1&\text{if }M[i,j]=1\\ S[i,j]&\text{otherwise}\end{cases}$ (10)
the dilation of the mask $\mathbf{M_{src}}{}$ using the structuring element K, where denotes the dilation operation and G is the Gaussian kernel. The asterisk *** denotes the convolution operation. $S^{\prime}$ is the final soft mask.
Gradual Boundary Transition: This involves generating a sequence of arrays where the value of 1 does not appear immediately but increases incrementally from 0 to 1. This can be achieved by interpolating between 0 and 1 across the sequence of arrays.
$\mathbf{M_{src}}{}(x,y)=\begin{cases}\mathbf{M_{src}}{}(x,y)\cdot\frac{t}{K},&\text{if }t\leq K\text{ and }\mathbf{M_{src}}{}(x,y)=1\\ \mathbf{M_{src}}{}(x,y),&\text{otherwise}\end{cases}$ (11)
In this equation, the value of $\mathbf{M_{src}}{}(x,y)$ is assumed to be 1 in the center area and 0 elsewhere. For the central region, the value linearly increases from 0 to 1 over the first $K$ steps. For the rest of the mask, the original value $\mathbf{M_{src}}{}(x,y)$ remains unchanged.
Several prevalent backbone diffusion models, including Stable Diffusion 2.1, are restricted to processing images in a square format. Resizing images to fit a square dimension can lead to substantial content distortion, adversely affecting the editing outcomes. Nevertheless, our findings demonstrate that our method exhibits a remarkable capacity for adaptation, allowing it to process images of any aspect ratio without compromise. As documented in this paper, we present all images in various ratios.
5 Evaluation
Figure 4: Qualitative comparison with different baselines. Note that those baseline methods were already equipped with some components of SwapAnything for precise control of the swapping region. Please check Sec. 5.1 for details.
5.1 Implementation Details
Here we introduce implementation details. In our paper, we used Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model. DreamBooth (Ruiz et al., 2023) is used to convert the concept into textual space. We used null-text inversion (Mokady et al., 2023) based on DDIM inversion (Song et al., 2020) to boosts the accuracy and reliability of the inversion.
For object mask, we first detect the object with Grounding DINO (Liu et al., 2023) and then extract the mask using Segment Anything (Kirillov et al., 2023). For the targeting variable swapping process, we do 30 for latent image feature $z$, 20 steps for cross-attention map, 25 for the self-attention map, and 10 for the self-attention output, We conduct swapping in all U-Net layer. There is no additional operation for single-object, partial object, cross domain swapping. Multi-object swapping is achieved by conducting swapping operation on the previous swapped image.
Dataset. We conducted experiments on both human and non-human objects. For human swapping, we collect celebrities from internet searches. We employed the search prompt: “a photo of $<target>$”, where $<target>$ is the celebrity name. We collected images of 15 celebrities for the concept learning process. We also collected 500 images containing 1 or more people as the source images. For non-human object, we include DreamEdit (Li et al., 2023b) dataset and more concepts and its corresponding source images from Google search. In total, we aggregated 1,000 images. We will also release the collected dataset.
Baselines setting. Photoswap(Gu et al., 2023a), P2P(Hertz et al., 2022), PnP(Tumanyan et al., 2023), MasaCtrl(Cao et al., 2023) are attention variable based image editing methods, which are also compatible with our proposed Masked Latent Blending in Sec. 4.1 and Location Adaptation in Sec. 4.2. Therefore we boost their performance with those additional components. Without those, their performance would be much worse. Please see appendix for comparisons with their original methods. Besides, we use external mask to help the inpainting process in DreamEdit. SwapAnything is also compared with BlipDiffusion (Li et al., 2023a). Photoswap, P2P, PnP, and MasaCtrl, DreamEdit were equipped with the same DreamBooth model to grasp the new concept. Note that this would also indirectly include comparison with CustomEdit (Choi et al., 2023), since it also achieved personalized object swapping via equiping P2P with concept learning. CopyPaste involves directly transplanting the concept object in the concept image into the source object’s position.
Table 1: Human evaluation results. We show the human preference between results generated by our method and the baseline methods. SS means Object Swapping, BP means Background Preservation, and OQ means Overall Quality. SG means Object Gesture. For the baseline methods, PS means Photoswap; MC means MasaCtrl; BP means BlipDiffusion; DE means DreamEdit; CP means CopyPaste.
Ours PS Tie Ours P2P Tie Ours PnP Tie Ours MC Tie Ours BP Tie Ours DE Tie Ours CP Tie
SS 52.3 12.5 35.2 55.3 17.9 26.8 52.3 29.9 11.5 64.3 20.0 15.7 55.3 16.5 28.2 55.1 20.1 24.8 60.1 17.3 22.6
SG 44.5 34.0 21.5 49.5 32.0 18.5 55.5 33.0 11.5 65.3 18.8 15.9 41.5 39.6 18.9 44.3 18.8 36.9 54.3 20.5 25.2
BP 41.5 32.0 26.5 44.5 28.9 26.6 50.2 20.9 28.9 55.2 23.1 21.7 40.0 27.5 32.5 44.0 18.9 37.1 54.1 13.1 32.8
OQ 49.3 27.8 22.9 55.0 27.2 17.8 52.5 22.9 24.6 59.4 20.1 20.5 48.1 25.3 26.6 53.1 19.9 27.0 71.1 10.5 18.4
5.2 Single-object Swapping
Table 2: Automatic evaluation results. SwapAnything outperforms all other methods across all metrics.
Ours Photoswap P2P PnP MasaCtrl BlipDiffusion DreamEdit CopyPaste
$DINO_{fore}$ 0.61 0.55 0.47 0.49 0.29 0.44 0.52 0.56
$CLIP_{fore}$ 0.79 0.53 0.71 0.73 0.46 0.54 0.61 0.75
$DINO_{back}$ 0.79 0.75 0.68 0.64 0.71 0.71 0.76 0.77
$CLIP_{back}$ 0.89 0.86 0.75 0.70 0.67 0.76 0.82 0.79
We consider human evaluation to be the main quantitative performance measurement. A successful swap should keep the non-object area unchanged, change the object identity to target, and keep the gesture the same as the source object. As in Tab. 1, our model consistently outperforms baselines across all metrics. Fig. 4 shows the qualitative comparison for both human and non-human images. Thanks to the addition of our targeting variable swapping and location adaptation, all attention-manipulation based baselines also achieved perfect background preservation and some level of localized swapping result. SwapAnything yield a much better appearance adaptation result. For example, in the first row, the human facial expression and the direction of the eyes are fully aligned with the source image.
We also conducted automatic evaluation. Following Ruiz et al. (2023); Li et al. (2023b); Gu et al. (2023a), we employ both DINO and CLIP-I as tools to evaluate the quality of the images generated. These two metrics serve as complementary indicators to the results obtained from human evaluations. As in Table 2, SwapAnything outperforms all other baselines in terms of both subject identity swapping and background preservation, which is consistent with the results of human evaluation.
5.3 Multi-object Swapping
As is shown in Fig. 5, multi-object swapping is simply achieved via repeating single-object swapping, which further highlights its versatility and efficiency. Multi-object swapping is a natural outcome of our targeted variable swapping, whereas previous methods struggle to achieve satisfactory results. Without perfect context pixel preservation, the unwanted image modification would accumulate as the swapping continues. Please check the appendix for more results and comparison with baselines.
Figure 5: Multi-object swapping results of SwapAnything. Given different source images, we swap the identity from the same concept object. Note that the red circle means the target object to be replaced. The same color means a pair of concept and target for object swapping.
5.4 Partial Object Swapping
As is shown in Fig. 6, SwapAnything achieved a great performance on swap a part of a whole object, even when the targeting area is very small. Meanwhile, all other baselines failed to achieve such results. Please check appendix for more results.
Figure 6: Results on partial object swapping. SwapAnything can swap partial object that is tightly connected with other parts and adapt seamlessly to the source image. The second row is the zoom-in images of the swapping part in the first row.
5.5 Cross-domain Swapping
Fig. 7 demonstrates that SwapAnything can adeptly handle a range of stylized source images, successfully adapting concept objects to match the desired style within the source image while seamlessly transferring identity into the generated images. For instance, in the fourth row, the source image is an oil painting photo, yet SwapAnything skillfully generates the same painting style featuring personalities like “Justin Bieber” and “Jackie Chan”. Notably, all the concept images of “Justin Bieber” and “Jackie Chan” are regular, unstyled photos, underscoring the model’s ability to blend different styles and identities effectively. Besides swapping tasks listed in Sec. 5, SwapAnything is also capable of other tasks such as swapping background. Please refer to the appendix for more results.
Figure 7: Results on cross-domain object swapping. With a variety of source images, including graphic, black-white photos, and oil paintings, the framework seamlessly integrates concept objects taken from regular internet images into these diverse source images.
5.6 Text-based Swapping
As shown in Fig. 8, besides personalized swapping, our method can also do text-based swapping, swapping an object in the source image with another described in text. This can be achieved by simply replacing the personalized concept token *** with a text prompt, e.g., “A photo of $new\_obj$”.
Figure 8: Results on text-based swapping. Besides swapping from concept image, SwapAnything is also capable of swapping an object described in text into the image.
5.7 Object Insertion
SwapAnything is a general framework and is also capable of object insertion. With the same process as single-object swapping, we could insert and adapt a concept into background pixels, while preserving the composition and style of the source image. In Fig. 9, we insert a puppy and a butterfly into The Starry Night from Vincent van Gogh.
Figure 9: Results on object insertion. SwapAnything can insert and adapt an object into a certain location of an image.
5.8 Ablation Study
Fig. 10 show the effect of the components in SwapAnything. From the left part, we see that without latent feature swap, even with a mask and attention variable swap, the context pixel such as clothes is still changed. Both latent feature and attention variable has effect of information preservation when compared with the result of no swap. The effect of adaptation and mask is presented on the right part. With style adaptation, the visual texture is closer to the source image. Without scale adaptation, the face shape is not well aligned and artifacts appear in the neck part. Without content adaptation, artifacts such as a hand touching the chin appear on the neck in the image. When without any adaptation, the generated image is much less connected the source image regarding the swapping area. When without mask, both background and targeting area are changed, which leads to a different image. Also, without the swapping and the adaptation proposed in SwapAnything, the edited image would have stong visual distortion after reshape to the size of source image, so we present those two images with model’s output size.
Figure 10: Ablation study. The left part shows the effect of swapping, and the right part shows the effect of adaptation and mask. Please refer to the appendix for more results.
5.9 Ethics Discussion
Existing text-to-image diffusion models (Ruiz et al., 2023; Rombach et al., 2022) can potentially exhibit biases reflective of those inherent in the training data. Given that these models are trained on vast text and image datasets, they might inadvertently learn and perpetuate biases. However, SwapAnything aims to mitigate such biases by incorporating style and content adaptations, as detailed in Sec. 4.2. Nevertheless, to ensure authenticity, we recommend using SwapAnything with objects of the same gender and similar racial background. Moreover, our intention for SwapAnything is to serve humanitarian purposes, such as creating group photos with deceased family members. We strictly prohibit the creation of harmful content.
6 Conclusion
In conclusion, SwapAnything represents a notable breakthrough in the realm of object swapping. Swapping latent features and attention variables in the diffusion model ensures the retention of crucial information within the generated image. Through a targeted manipulation, we achieved a perfect background preservation. Additionally, we have introduced a sophisticated appearance adaptation process designed to seamlessly integrate the concept into the context of the source image. Consequently, SwapAnything is equipped to handle a diverse array of object swapping challenges. In the future, we plan to extend our framework to 3D/video personalized object swapping tasks.
Appendix A Variable Swapping Details
We use Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model. DreamBooth (Ruiz et al., 2023) is used to convert the concept into textual space. The learning rate for this process is set at 1e-6, and we use the Adawm optimizer for 800 steps. The U-net and the text encoder are fine-tuned during this process, typically taking about 2 minutes on a machine equipped with 8 A100 GPUs. The target prompt is essentially the source prompt with a swap in object tokens to introduce a new concept.
For area mask smooth, we first enlarge the masked areas using a dilation operation with an elliptical kernel, which can be adjusted in size. After dilation, the mask edges are smoothed using a Gaussian blur, creating a gradient effect at the boundaries. For the smooth over diffusion step, we linearly increase the mask rate from 0 to 1 during the first 30 steps. For a better understanding, we mark the masked area using a circle in most figures in this paper.
Appendix B Adaptation Details
Style Adaptation. This operation adjusts the mean and variance of content image features to match those of the style features, facilitating the transfer of artistic styles onto content images. The AdaIN technique is renowned for its efficiency and flexibility, making it a go-to choice for real-time style transfer and artistic image manipulation. Building on this, we introduce Masked-AdaIN. Unlike traditional AdaIN which applies style alignment across the entire image, Masked-AdaIN focuses this alignment only on a specific target area. In this approach, mean and variance calculations are exclusively performed on the designated masked area, allowing for more precise and localized style transfers.
Scale Adaptation. We adapt the scale of the object in latent space to the shape of the mask. The object shape is indicated in the cross-attention map at each diffusion step (Gu et al., 2023a; Hertz et al., 2022). $Shape(\mathbf{M_{src}}{})(k)$ means the attention map for object text token $k$, which is obtained through binary-like transformation to the attention map. We apply a threshold of 0.4 after using sigmoid to normalize the attention value between 0 and 1.
Content Adaptation. In the Linear Boundary Interpolation process, the structuring element $K$ is a predefined shape used in the dilation process to create the dilated image. The structuring element $K$ slides over the binary mask $\mathbf{M_{src}}{}$ and at each position. If at least one pixel under $K$ is 1, the pixel in the output image under the center of $K$ is set to 1. This operation typically results in the enlargement of the regions with 1s in the binary mask, effectively smoothing the boundary and filling small holes and gaps. The subsequent convolution with a Gaussian kernel $G$ further smooths the mask by averaging values in the vicinity of each point, thereby creating a gradient effect. The combination of dilation and Gaussian smoothing prepares the mask $S^{\prime}$ for linear boundary interpolation, where the sharp transitions are made gradual, and the final soft mask $S^{\prime}$ is obtained by selectively setting pixels to 1 based on the original mask and the smoothed values. In Gradual Boundary Transition, we set the transition step parameter as 30 to anneal $\mathbf{M_{src}}{}$ from 0 to the set value.
Appendix C More Qualitative Results
Figure 11: Comparison on single-object swapping with baselines in their original components. SS means Object Swapping, BP means Background Preservation, and OQ means Overall Quality. SG means Object Gesture. Please zoom in for a clear visual result.
Here we first show the comparison with baselines in their original setting on single-object swapping. On other more challenging tasks, we also show the results of Photoswap since it is the state-of-the-art method of subject swapping.
Single-object Swapping. Fig. 11 shows comparisons between SwapAnything and baselines. SwapAnything consistently outperforms other models in terms of background preservation, identity swapping, and overall quality. Note that there is also a huge performance gap between some baselines and their counterpart in Fig. 4 in the main paper, which further validates the efficacy of targeted variable swapping and location adaptation, which was applied to Photoswap, P2P, PnP, and MasaCtrl in Fig. 4 in the main paper.
Partial Object Swapping. As in Fig. 12, our method precisely swaps the cat head with a raccoon head harmoniously without influencing other pixels. Meanwhile, Photoswap swaps the whole body and modified the context pixels. When our proposed masked variable swapping is added, Photoswap achieves a better background preservation performance.
Cross-domain Swapping. SwapAnything is capable of swapping between styles and textual. In Fig. 12, a bear is adapted into a logo while keeping the gesture of the source object horse. Meanwhile, Photoswap fails to complete the challenging task. Also, when masked variable swapping is added, Photoswap achieves a better adaptation performance.
Figure 12: Comparison with Photoswap on partial object swapping and cross-domain swapping. The upper part shows SwapAnything could localize the swapping area while Photoswap inevitably modified the background. In the lower part, SwapAnything adapts a bear into the style of a logo, while Photoswap failed on this cross-domain swapping task.
Multi-object Swapping. Multi-object swapping is a big step after single-object swapping. Here in Fig. 13, we show two main obstacles we resolved. First, previous methods usually have a background modification such that continuous editing would accumulate unwanted distortion, which leads to a totally different image and fails the task of swapping. The second issue is that previous methods are usually designed for main subject swapping and do not pay attention to other objects. In this case, the objects in the following swapping steps could disappear in the previous swapping process.
Figure 13: Comparison between SwapAnything and Photoswap on multi-object swapping. Photoswap usually has a large modification on the context pixel and cannot precisely keep all objects so it fails to conduct continuous swapping.
Additional Ablation Study We present one more example to show the effect of various components in SwapAnything. This cross-domain swapping example is more challenging than the one in the main paper, which further shows the importance of each module.
Figure 14: Additional ablation study. The swapping area is the compass logo in the middle of the bag.
Appendix D More Qualitative Results
Tab. 3 shows results on human evaluation on both human and non-human images on top baselines. PS means Photoswap (Gu et al., 2023a); P2P means Prompt-to-Prompt (Hertz et al., 2022); PnP means Plug-and-Play (Tumanyan et al., 2023); DE means DreamEdit (Li et al., 2023b). We also conduct comparisons with another baseline PbE, Paint-by-Example (Yang et al., 2023a).
Table 3: User study results. The 2nd to 5th rows and 6th to 9th rows show results on human objects and non-human objects.
Ours PS Tie Ours P2P Tie Ours PnP Tie Ours DE Tie Ours PbE Tie
SS 59.0 10.0 31.0 52.7 18.2 24.1 58.8 29.2 12.0 53.4 16.5 30.1 62.1 12.0 28.0
SG 44.0 33.7 22.3 54.5 29.1 16.4 61.6 33.3 5.1 42.4 17.2 40.4 73.1 15.8 12.0
BP 45.4 32.2 22.4 49.9 26.9 23.2 49.7 22.0 28.3 43.8 18.0 38.2 42.1 31.3 27.0
OQ 47.3 24.3 28.4 58.4 23.3 18.3 51.6 31.1 17.3 47.5 26.5 26.0 52.1 21.9 30.0
SS 45.6 15.0 39.4 52.9 17.6 29.5 45.8 30.6 23.6 56.8 23.7 19.5 58.1 12.0 27.8
SG 45.0 34.3 20.7 44.5 34.9 20.6 49.4 32.7 17.9 46.2 20.4 33.4 69.3 15.0 14.8
BP 37.6 31.8 30.6 39.1 30.9 30.0 50.7 19.8 29.5 44.2 19.8 36.0 40.5 31.5 27.6
OQ 51.3 31.3 17.4 51.6 31.1 17.3 47.5 26.5 26.0 48.1 21.2 30.7 48.1 18.3 29.6
Appendix E Failure Cases
We highlight two common failure scenarios encountered in our experiments. Firstly, although our model adeptly preserves contextual pixels, it struggles with accurately reconstructing non-target pixels within the mask area. This issue is illustrated in the first row of Figure 15, the position of ring in the source image is changed in the target image. The second challenge arises when dealing with subjects that exhibit a high degree of variability or freedom of movement. In such cases, as shown in the bottom row of the figure, accurately replicating the concept subject becomes difficult. To address this, we are considering the implementation of explicit alignment, which we aim to explore in our future work.
Figure 15: Examples of failure cases. The model sometimes struggles to keep the details inside the mask area and could fail if the object has a high degree of freedom.
Appendix F Human Evaluation Interface
Figure 16: The illustration of the user study interface.
Amazon Turker was presented with one reference image mainly containing the concept subject, one source image to be swapped, and two generated images from SwapAnything and a baseline.

### v3

@@
-SwapAnything: Enabling Arbitrary Object Swapping in Personalized Visual Editing
+https://swap-anything.github.io/
 Abstract
@@
 SwapAnything is adept at precise, arbitrary object replacement in a source image with a personalized reference, and achieves high-fidelity swapping results without influencing any context pixels.
-We demonstrate its general effect in single-object, multi-object, partial-object, and cross-domain swapping tasks.
+We demonstrate its general swapping effect in single-object, multi-object, partial-object, and cross-domain swapping tasks.
 1 Introduction
-In today’s digital age marked by the prolific creation and widespread sharing of personal content, generative models (Rombach et al., 2022; Crowson et al., 2022; Ding et al., 2021; Blattmann et al., 2022) have risen as a potent tool for self-expression, storytelling, and amplifying the impact of visual narratives.
-Among existing image editing methods (Meng et al., 2022; Hertz et al., 2022; Chen et al., 2023b) for content creation that empowered individuals to convey their creative instincts, weave captivating narratives into their visual stories, and enhance the quality of their visual representations, personalize content swapping (Gu et al., 2023a; Li et al., 2023b; Gu et al., 2023b), which targets at creating and compositing new images with user-specified visual concept, attracted significant interest due to its wide-ranging applications in E-commerce, entertainment, and professional editing.
+In today’s digital age marked by the prolific creation and widespread sharing of personal content, generative models [32, 7, 9, 2] have risen as a potent tool for self-expression, storytelling, and amplifying the impact of visual narratives.
+Among existing image editing methods [27, 15, 5] for content creation that empowered individuals to convey their creative instincts, weave captivating narratives into their visual stories, and enhance the quality of their visual representations, personalize content swapping [13, 22, 14], which targets at creating and compositing new images with user-specified visual concept, attracted significant interest due to its wide-ranging applications in E-commerce, entertainment, and professional editing.
 Achieving arbitrary personalized content swapping necessitates a deep understanding of the visual concept inherent to both the original and replacement subjects.
@@
 Existing works often fall short of addressing these challenges.
-Most of existing research (Ruiz et al., 2023; Shi et al., 2023; Chen et al., 2023b; Jia et al., 2023; Wang et al., 2024) are focused on personalized image synthesis, which seeks to create new images with personalized content.
+Most of existing research [34, 36, 5, 18, 42] are focused on personalized image synthesis, which seeks to create new images with personalized content.
 Although these approaches can synthesize high-fidelity content, they cannot edit or replace the visual content in an existing image.
-Li et al.
-(2023b); Cao et al.
-(2023); Meng et al.
-(2022); Avrahami et al.
-(2023) have tried to remove and regenerate the object via masks, they often struggle to adapt the new concept into the image.
-Recently studies (Patashnik et al., 2023; Gu et al., 2023a; Tumanyan et al., 2023) focus on object swapping and replacement by tweaking intermediate variables affecting the object’s features.
+[22, 3, 27, 1] have tried to remove and regenerate the object via masks, they often struggle to adapt the new concept into the image.
+Recently studies [31, 13, 40] focus on object swapping and replacement by tweaking intermediate variables affecting the object’s features.
 However, this approach lacks the precision needed for localized object swapping, resulting in limited visual qualities.
@@
 2.1 Text-guided Image Editing
-With recent progress in diffusion model, text-guided Image editing has been largely explored (Avrahami et al., 2023).
+With recent progress in diffusion model, text-guided Image editing has been largely explored [1].
 Image editing driven by text aims to alter an existing image following the textual guidelines provided, while ensuring certain elements or features of the original image remain unchanged.
-Initial efforts using GAN models (Karras et al., 2019) were restricted to specific object domains.
-However, diffusion-oriented techniques (Zhang et al., 2023; Nichol et al., 2022; Feng et al., 2023) have surpassed this limitation, enabling text-driven image modifications.
+Initial efforts using GAN models [19] were restricted to specific object domains.
+However, diffusion-oriented techniques [46, 29, 11] have surpassed this limitation, enabling text-driven image modifications.
 While these techniques produce captivating outcomes, many grapple with localized edits.
-As a result, manual masks (Meng et al., 2022; Zeng et al., 2023; Meng et al., 2022) are often help to delineate the editing zones.
-Though employing cross-attention (Hertz et al., 2022) or spatial attributes (Tumanyan et al., 2023) can achieve localized edits, they face challenges with flexible changes (like altering poses) and preserving the initial image composition.
+As a result, manual masks [27, 45, 27] are often help to delineate the editing zones.
+Though employing cross-attention [15] or spatial attributes [40] can achieve localized edits, they face challenges with flexible changes (like altering poses) and preserving the initial image composition.
 2.2 Object Driven Image Editing
-In object-driven image generation, the object could be inverted into the texual space so to represented by a new token such as “***”  (Gal et al., 2023; Ruiz et al., 2023; Chen et al., 2023a; Shi et al., 2023; Tewel et al., 2023).
+In object-driven image generation, the object could be inverted into the texual space so to represented by a new token such as “*”  [12, 34, 4, 36, 39].
 Image editing guided by exemplars spans a vast array of uses.
-Much of the research in this area (Wang et al., 2019; Huang et al., 2018; Zhou et al., 2021) falls under the umbrella of example-based image transformation tasks.
-These tasks leverage various sources of information, from stylized images (Liu et al., 2021; Deng et al., 2022; Zhang et al., 2022) and layouts (Yang et al., 2023b; Li et al., 2023c; Jahn et al., 2021), to skeletons (Li et al., 2023c) and sketches or outlines (Seo et al., 2023).
-Given the versatility of stylized images, the art of image style adaptation (Liao et al., 2017; Zhang et al., 2020) has garnered significant interest.
+Much of the research in this area [41, 16, 49] falls under the umbrella of example-based image transformation tasks.
+These tasks leverage various sources of information, from stylized images [26, 8, 48] and layouts [44, 23, 17], to skeletons [23] and sketches or outlines [35].
+Given the versatility of stylized images, the art of image style adaptation [24, 47] has garnered significant interest.
 These methods primarily focus on establishing a detailed match between the input and reference visuals but falter when it comes to localized alterations.
@@
 However, these tools often demand manual input, which can be challenging for users.
-A novel approach (Yang et al., 2023a) frames exemplar-driven image editing as a task of filling in gaps, maintaining the context while transferring semantic elements from the reference to the original image.
-While DreamEdit (Li et al., 2023b) employs a step-by-step gap-filling method for object substitution, it doesn’t effectively bridge the connection between the original and desired objects.
+A novel approach [43] frames exemplar-driven image editing as a task of filling in gaps, maintaining the context while transferring semantic elements from the reference to the original image.
+While DreamEdit [22] employs a step-by-step gap-filling method for object substitution, it doesn’t effectively bridge the connection between the original and desired objects.
 In contrast, our method ensures that crucial features, like body movements and facial expressions, stay consistent.
-Meanwhile, Photoswap (Gu et al., 2023a), CustomEdit (Choi et al., 2023) and BLIP-Diffusion (Li et al., 2023a) achieved personalized object swapping.
+Meanwhile, Photoswap [13], CustomEdit [6] and BLIP-Diffusion [21] achieved personalized object swapping.
 However, it does not keep the non-object pixel background intact.
@@
 Diffusion models predict and reverse the noise at each step in the diffusion sequence to arrive at the final denoised image.
-In our research, we employ the pre-trained text-to-image diffusion framework Stable Diffusion Rombach et al.
-(2022).
+In our research, we employ the pre-trained text-to-image diffusion framework Stable Diffusion [32].
 This model encodes images into a latent space and incrementally denoises the encoded latent representation.
-The Stable Diffusion model operates on a U-Net architecture (Ronneberger et al., 2015), where the latent representation ${\boldsymbol{z}}_{t-1}$ at any given step is derived from the text prompt $P$ and the previous latent state ${\boldsymbol{z}}_{t}$, as indicated by the following equation:
+The Stable Diffusion model operates on a U-Net architecture [33], where the latent representation ${\boldsymbol{z}}_{t-1}$ at any given step is derived from the text prompt $P$ and the previous latent state ${\boldsymbol{z}}_{t}$, as indicated by the following equation:
 ${\boldsymbol{z}}_{t-1}={\boldsymbol{\epsilon\theta}}({\boldsymbol{z}}_{t},P,t)$ (1)
@@
 where $A$ is the cross-attention map. In this work, we study the swapping of $A$, $M$, $\phi$ and $z$.
-Figure 2: Overview of SwapAnything on swapping a object from a source image ($I_{src}$) into a personalized concept (<*><{*}>< * >) to get the target image ($I_{target}$).
+Figure 2: Overview of SwapAnything on swapping a object from a source image ($I_{src}$) into a personalized concept (<><{*}>< ∗ >) to get the target image ($I_{target}$).
 The personalized concept is first converted into textual space to be treated as concept appearance.
@@
 4.1 Targeted Variable Swapping
-Intermediate variables in the U-Net of a diffusion model have been proven informative about the content of the generated image (Gu et al., 2023a; Hertz et al., 2022; Cao et al., 2023; Tumanyan et al., 2023).
+Intermediate variables in the U-Net of a diffusion model have been proven informative about the content of the generated image [13, 15, 3, 40].
 They usually focus on the study of variables inside of U-net structure such as attention map, and attention output, while the output of U-Net at each diffusion step, i.e., latent image feature ${\boldsymbol{z}}$ is not widely explored before.
@@
 Swapping those facilitates the alignment of the latent features between the source and target images before the partial-swapping between them.
-In short, all the variables mentioned above in both the source image and target image generation process could be resized into the shape of the mask, where the mask can be utilized for the swapping process.
+In short, all the variables mentioned above in both the source image and target image generation process could be resized into the shape of the mask, where the mask can be utilized for the swapping process to get target variable:
 $V_{target}=g(f(V_{src})*(1-\mathbf{M_{src}}{})+f(V_{target})*\mathbf{M_{src}}{})$ (4)
@@
 However, we find that a changing mask usually confuses the model and leads to a less optimal performance.
-Therefore, while using the same high-quality mask through the diffusion process.
-We find that the mask could either be extracted from the source image directly using an off-the-shelf model or from the generation process as in Simsar et al.
-(2023); Patashnik et al.
-(2023).
+Therefore, we use the same high-quality mask through the diffusion process.
+We find that the mask could either be extracted from the source image directly using an off-the-shelf model or from the generation process as in [37, 31].
 Please check the Appendix for more details.
@@
 Therefore, solely generating the foreground information via the text prompt might lead to style inconsistency.
-Recently, Karras et al.
-(2019); Park et al.
-(2019) found that adding such normalization layers can help improve the conditional image generation quality because such activation functions modulation.
-Unlike from them, in our work, we employ the AdaIN (Adaptive Instance Normalization) to modulate the swapping features with spatial constraints.
+Recently, [19, 30] found that adding such normalization layers can help improve the conditional image generation quality because such activation functions modulation.
+Unlike them, we employ the AdaIN (Adaptive Instance Normalization) to modulate the swapping features with spatial constraints.
 We follow Eq.
@@
 Guidance from an external classifier in the inference process of Diffusion models could influence the diffusion noise to control the generated image.
-Epstein et al.
-(2023) shows that the guidance can also be used on the attention map to control the generation.
+[10] shows that the guidance can also be used on the attention map to control the generation.
 Similarly, we adapt the mask guidance Eq.
@@
 As with classifier guidance, we scale by $\sigma_{t}$ to convert the score function to a prediction of $\mathbf{\epsilon}_{t}$.
-$Shape(\mathbf{M_{src}}{})$ denotes the object shape as identified in the attention layer.
-Here the energy function $g$ is set as $\|\mathbf{M_{src}}{}-Shape(\mathbf{M_{src}}{})\|_{1}$ to calculate the shape difference between the original object mask and the extracted shape in the attention layer, which indicates the deviation between the ideal shape and shape during the diffusion process.
+$Shape(\mathbf{M_{src}}{})\left(k\right)$ denotes the object shape as identified in the cross-attention layer.
+Here the energy function $g$ is set as $\|\mathbf{M_{src}}{}-Shape(\mathbf{M_{src}}{})\left(k\right)\|_{1}$ to calculate the shape difference between the original object mask and the extracted shape of object token $k$ in the attention layer, which indicates the deviation between the ideal shape and shape during the diffusion process.
 4.2.4 Content Adaptation
@@
 The dilation of the mask $\mathbf{M_{src}}{}$ using the structuring element K, where denotes the dilation operation and G is the Gaussian kernel.
-The asterisk *** denotes the convolution operation.
+The asterisk * denotes the convolution operation.
 $S^{\prime}$ is the final soft mask.
@@
 In our paper, we used Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model.
-DreamBooth (Ruiz et al., 2023) is used to convert the concept into textual space.
-We used null-text inversion (Mokady et al., 2023) based on DDIM inversion (Song et al., 2020) to boosts the accuracy and reliability of the inversion.
-For object mask, we first detect the object with Grounding DINO (Liu et al., 2023) and then extract the mask using Segment Anything (Kirillov et al., 2023).
-For the targeting variable swapping process, we do 30 for latent image feature $z$, 20 steps for cross-attention map, 25 for the self-attention map, and 10 for the self-attention output, We conduct swapping in all U-Net layer.
+DreamBooth [34] is used to convert the concept into textual space.
+We used null-text inversion [28] based on DDIM inversion [38] to boost the accuracy.
 There is no additional operation for single-object, partial object, cross domain swapping.
 Multi-object swapping is achieved by conducting swapping operation on the previous swapped image.
-Dataset.
-We conducted experiments on both human and non-human objects.
-For human swapping, we collect celebrities from internet searches.
-We employed the search prompt: “a photo of $<target>$”, where $<target>$ is the celebrity name.
-We collected images of 15 celebrities for the concept learning process.
-We also collected 500 images containing 1 or more people as the source images.
-For non-human object, we include DreamEdit (Li et al., 2023b) dataset and more concepts and its corresponding source images from Google search.
-In total, we aggregated 1,000 images.
-We will also release the collected dataset.
+Please check Appendix for evaluation dataset details.
 Baselines setting.
-Photoswap(Gu et al., 2023a), P2P(Hertz et al., 2022), PnP(Tumanyan et al., 2023), MasaCtrl(Cao et al., 2023) are attention variable based image editing methods, which are also compatible with our proposed Masked Latent Blending in Sec.
+Photoswap[13], P2P[15], PnP[40], MasaCtrl[3] are attention variable based image editing methods, which are also compatible with our proposed Masked Latent Blending in Sec.
 4.1 and Location Adaptation in Sec.
@@
 Therefore we boost their performance with those additional components.
-Without those, their performance would be much worse.
-Please see appendix for comparisons with their original methods.
-Besides, we use external mask to help the inpainting process in DreamEdit.
-SwapAnything is also compared with BlipDiffusion (Li et al., 2023a).
-Photoswap, P2P, PnP, and MasaCtrl, DreamEdit were equipped with the same DreamBooth model to grasp the new concept.
-Note that this would also indirectly include comparison with CustomEdit (Choi et al., 2023), since it also achieved personalized object swapping via equiping P2P with concept learning.
-CopyPaste involves directly transplanting the concept object in the concept image into the source object’s position.
+Otherwise, their performance would be much worse.
+Please see appendix for comparisons with their original methods, and the implementation details.
+Also, please check appendix for its performance on object insertion.
 Table 1: Human evaluation results.
@@
 5.2 Single-object Swapping
-Table 2: Automatic evaluation results. SwapAnything outperforms all other methods across all metrics.
-Ours Photoswap P2P PnP MasaCtrl BlipDiffusion DreamEdit CopyPaste
-$DINO_{fore}$ 0.61 0.55 0.47 0.49 0.29 0.44 0.52 0.56
-$CLIP_{fore}$ 0.79 0.53 0.71 0.73 0.46 0.54 0.61 0.75
-$DINO_{back}$ 0.79 0.75 0.68 0.64 0.71 0.71 0.76 0.77
-$CLIP_{back}$ 0.89 0.86 0.75 0.70 0.67 0.76 0.82 0.79
 We consider human evaluation to be the main quantitative performance measurement.
@@
 SwapAnything yield a much better appearance adaptation result.
-For example, in the first row, the human facial expression and the direction of the eyes are fully aligned with the source image.
-We also conducted automatic evaluation.
-Following Ruiz et al.
-(2023); Li et al.
-(2023b); Gu et al.
-(2023a), we employ both DINO and CLIP-I as tools to evaluate the quality of the images generated.
-These two metrics serve as complementary indicators to the results obtained from human evaluations.
-As in Table 2, SwapAnything outperforms all other baselines in terms of both subject identity swapping and background preservation, which is consistent with the results of human evaluation.
 5.3 Multi-object Swapping
@@
 Figure 5: Multi-object swapping results of SwapAnything.
-Given different source images, we swap the identity from the same concept object.
+Our method could easily swap multiple objects via swapping one object at a time.
 Note that the red circle means the target object to be replaced.
@@
 7 demonstrates that SwapAnything can adeptly handle a range of stylized source images, successfully adapting concept objects to match the desired style within the source image while seamlessly transferring identity into the generated images.
-For instance, in the fourth row, the source image is an oil painting photo, yet SwapAnything skillfully generates the same painting style featuring personalities like “Justin Bieber” and “Jackie Chan”.
-Notably, all the concept images of “Justin Bieber” and “Jackie Chan” are regular, unstyled photos, underscoring the model’s ability to blend different styles and identities effectively.
-Besides swapping tasks listed in Sec.
-5, SwapAnything is also capable of other tasks such as swapping background.
-Please refer to the appendix for more results.
+For instance, when the source image is a photo of a certain style, yet SwapAnything skillfully generates the same painting style featuring personalities like “Charles Darwin” and “J.
+Robert Oppenheimer”.
+Notably, all the concept images are regular, unstyled photos, underscoring the model’s ability to blend different styles and identities effectively.
 Figure 7: Results on cross-domain object swapping.
-With a variety of source images, including graphic, black-white photos, and oil paintings, the framework seamlessly integrates concept objects taken from regular internet images into these diverse source images.
+With a variety of source images, including graphic, black-white photos, and oil paintings, the framework seamlessly integrates concept objects taken from regular images into these diverse source images.
 5.6 Text-based Swapping
@@
 8, besides personalized swapping, our method can also do text-based swapping, swapping an object in the source image with another described in text.
-This can be achieved by simply replacing the personalized concept token *** with a text prompt, e.g., “A photo of $new\_obj$”.
+This can be achieved by simply replacing the personalized concept token * with a text prompt, e.g., “A photo of $new\_obj$”.
 Figure 8: Results on text-based swapping. Besides swapping from concept image, SwapAnything is also capable of swapping an object described in text into the image.
-5.7 Object Insertion
-SwapAnything is a general framework and is also capable of object insertion.
-With the same process as single-object swapping, we could insert and adapt a concept into background pixels, while preserving the composition and style of the source image.
-In Fig.
-9, we insert a puppy and a butterfly into The Starry Night from Vincent van Gogh.
-Figure 9: Results on object insertion. SwapAnything can insert and adapt an object into a certain location of an image.
 5.7 Ablation Study
@@
 9 show the effect of the components in SwapAnything.
-From the left part, we see that without latent feature swap, even with a mask and attention variable swap, the context pixel such as clothes is still changed.
+From the left part, we see that without latent feature swap, even with a mask and attention variable swap, the context pixel is still changed.
 Both latent feature and attention variable has effect of information preservation when compared with the result of no swap.
@@
 With style adaptation, the visual texture is closer to the source image.
-Without scale adaptation, the face shape is not well aligned and artifacts appear in the neck part.
-Without content adaptation, artifacts such as a hand touching the chin appear on the neck in the image.
+Without scale adaptation, the shape is not well aligned and artifacts appear in the neck part.
 When without any adaptation, the generated image is much less connected the source image regarding the swapping area.
 When without mask, both background and targeting area are changed, which leads to a different image.
-Also, without the swapping and the adaptation proposed in SwapAnything, the edited image would have stong visual distortion after reshape to the size of source image, so we present those two images with model’s output size.
-Figure 10: Ablation study. The left part shows the effect of swapping, and the right part shows the effect of adaptation and mask. Please refer to the appendix for more results.
+Figure 9: Ablation study. The left part shows the effect of swapping, and the right part shows the effect of adaptation and mask.
 5.8 Ethics Discussion
-Existing text-to-image diffusion models (Ruiz et al., 2023; Rombach et al., 2022) can potentially exhibit biases reflective of those inherent in the training data.
-Given that these models are trained on vast text and image datasets, they might inadvertently learn and perpetuate biases.
-However, SwapAnything aims to mitigate such biases by incorporating style and content adaptations, as detailed in Sec.
-4.2.
-Nevertheless, to ensure authenticity, we recommend using SwapAnything with objects of the same gender and similar racial background.
-Moreover, our intention for SwapAnything is to serve humanitarian purposes, such as creating group photos with deceased family members.
-We strictly prohibit the creation of harmful content.
+While SwapAnything achieves impressive performance across various visual tasks, it also raises potential ethical concerns.
+Like other advanced image editing technologies, SwapAnything could be misused to create deceptive content.
+To mitigate such risks, implementing technologies like digital watermarking can help track modifications and authenticate the sources of images, assisting users in distinguishing between genuine and altered content.
+Moreover, advocating for clear guidelines and regulations that govern the ethical use of image manipulation technologies is essential.
+These measures ensure responsible development and use, fostering a balance between innovation and ethical considerations.
+Here, we emphasize the importance of applications involving human face swapping.
+We strongly discourage using our methods on human faces without consent, and we are committed to preventing misinformation and harm to any individual or community.
+We enforce strict consent protocols in related tasks, requiring explicit permission from individuals before their images are used, especially in sensitive applications like face-swapping.
 6 Conclusion
-In conclusion, SwapAnything represents a notable breakthrough in the realm of object swapping.
+SwapAnything represents a notable breakthrough in the realm of object swapping.
 Swapping latent features and attention variables in the diffusion model ensures the retention of crucial information within the generated image.
@@
 In the future, we plan to extend our framework to 3D/video personalized object swapping tasks.
-Appendix A Variable Swapping Details
+Appendix 0.A Object Insertion
+SwapAnything is a general framework and is also capable of object insertion.
+With the same process as single-object swapping, we could insert and adapt a concept into background pixels, while preserving the composition and style of the source image.
+In Fig.
+10, we insert a puppy and a butterfly into The Starry Night from Vincent van Gogh.
+Figure 10: Results on object insertion. SwapAnything can insert and adapt an object into a certain location of an image.
+Appendix 0.B Variable Swapping Details
 We use Stable Diffusion 2.1 as the pre-trained text-to-image diffusion model.
-DreamBooth (Ruiz et al., 2023) is used to convert the concept into textual space.
+DreamBooth [34] is used to convert the concept into textual space.
 The learning rate for this process is set at 1e-6, and we use the Adawm optimizer for 800 steps.
@@
 The target prompt is essentially the source prompt with a swap in object tokens to introduce a new concept.
+For object mask, we first detect the object with Grounding DINO [25] and then extract the mask using Segment Anything [20].
+For the targeting variable swapping, we do 30 for latent image feature $z$, 20 steps for cross-attention map, 25 for the self-attention map, and 10 for the self-attention output, we conduct swapping in all U-Net layer.
 For area mask smooth, we first enlarge the masked areas using a dilation operation with an elliptical kernel, which can be adjusted in size.
@@
 For a better understanding, we mark the masked area using a circle in most figures in this paper.
-Appendix B Adaptation Details
+Appendix 0.C Adaptation Details
 Style Adaptation.
@@
 We adapt the scale of the object in latent space to the shape of the mask.
-The object shape is indicated in the cross-attention map at each diffusion step (Gu et al., 2023a; Hertz et al., 2022).
+The object shape is indicated in the cross-attention map at each diffusion step [13, 15].
 $Shape(\mathbf{M_{src}}{})(k)$ means the attention map for object text token $k$, which is obtained through binary-like transformation to the attention map.
@@
 In Gradual Boundary Transition, we set the transition step parameter as 30 to anneal $\mathbf{M_{src}}{}$ from 0 to the set value.
-Appendix C More Qualitative Results
-Figure 11: Comparison on single-object swapping with baselines in their original components.
-SS means Object Swapping, BP means Background Preservation, and OQ means Overall Quality.
-SG means Object Gesture.
-Please zoom in for a clear visual result.
+Appendix 0.D Dataset
+We conducted experiments on both human and non-human objects.
+For human swapping, we collect 50 faces as concepts.
+We also collected 500 images containing 1 or more people as the source images.
+For non-human object, we include DreamEdit [22] dataset and more concepts and its corresponding source images.
+In total, we aggregated 1,000 images.
+Appendix 0.E Human Evaluation Interface
+Figure 11: The illustration of the user study interface.
+Amazon Turker was presented with one reference image mainly containing the concept subject, one source image to be swapped, and two generated images from SwapAnything and a baseline.
+Appendix 0.F More Qualitative Results
+Figure 12: Comparison on single-object swapping with baselines in their original components. Please zoom in for a clear visual result.
 Here we first show the comparison with baselines in their original setting on single-object swapping.
 On other more challenging tasks, we also show the results of Photoswap since it is the state-of-the-art method of subject swapping.
+For the implementation details, we use external mask to help the inpainting process in DreamEdit.
+SwapAnything is also compared with BlipDiffusion [21].
+Photoswap, P2P, PnP, and MasaCtrl, DreamEdit were equipped with the same DreamBooth model to grasp the new concept.
+Note that this would also indirectly include comparison with CustomEdit [6], since it also achieved personalized object swapping via equiping P2P with concept learning.
+CopyPaste involves directly transplanting the concept object in the concept image into the source object’s position.
 Single-object Swapping.
@@
 As in Fig.
-12, our method precisely swaps the cat head with a raccoon head harmoniously without influencing other pixels.
+13, our method precisely swaps the cat head with a raccoon head harmoniously without influencing other pixels.
 Meanwhile, Photoswap swaps the whole body and modified the context pixels.
@@
 In Fig.
-12, a bear is adapted into a logo while keeping the gesture of the source object horse.
+13, a bear is adapted into a logo while keeping the gesture of the source object horse.
 Meanwhile, Photoswap fails to complete the challenging task.
@@
 Multi-object swapping is a big step after single-object swapping.
-Here in Fig.
-13, we show two main obstacles we resolved.
 First, previous methods usually have a background modification such that continuous editing would accumulate unwanted distortion, which leads to a totally different image and fails the task of swapping.
@@
 In this case, the objects in the following swapping steps could disappear in the previous swapping process.
-Figure 13: Comparison between SwapAnything and Photoswap on multi-object swapping.
-Photoswap usually has a large modification on the context pixel and cannot precisely keep all objects so it fails to conduct continuous swapping.
-Additional Ablation Study We present one more example to show the effect of various components in SwapAnything.
-This cross-domain swapping example is more challenging than the one in the main paper, which further shows the importance of each module.
-Figure 14: Additional ablation study. The swapping area is the compass logo in the middle of the bag.
-Appendix D More Qualitative Results
+Appendix 0.G More Quantitative Results
+Table 2: Automatic evaluation results. SwapAnything outperforms all other methods across all metrics.
+Ours Photoswap P2P PnP MasaCtrl BlipDiffusion DreamEdit CopyPaste
+$DINO_{fore}$ 0.61 0.55 0.47 0.49 0.29 0.44 0.52 0.56
+$CLIP_{fore}$ 0.79 0.53 0.71 0.73 0.46 0.54 0.61 0.75
+$DINO_{back}$ 0.79 0.75 0.68 0.64 0.71 0.71 0.76 0.77
+$CLIP_{back}$ 0.89 0.86 0.75 0.70 0.67 0.76 0.82 0.79
+We also conducted automatic evaluation.
+Following [34, 22, 13], we employ both DINO and CLIP-I as tools to evaluate the quality of the images generated.
+These two metrics serve as complementary indicators to the results obtained from human evaluations.
+As in Tab.
+2, SwapAnything outperforms all other baselines in terms of both subject identity swapping and background preservation, which is consistent with the results of human evaluation.
 Tab.
 3 shows results on human evaluation on both human and non-human images on top baselines.
-PS means Photoswap (Gu et al., 2023a); P2P means Prompt-to-Prompt (Hertz et al., 2022); PnP means Plug-and-Play (Tumanyan et al., 2023); DE means DreamEdit (Li et al., 2023b).
-We also conduct comparisons with another baseline PbE, Paint-by-Example (Yang et al., 2023a).
+PS means Photoswap [13]; P2P means Prompt-to-Prompt [15]; PnP means Plug-and-Play [40]; DE means DreamEdit [22].
+We also conduct comparisons with another baseline PbE, Paint-by-Example [43].
 Table 3: User study results. The 2nd to 5th rows and 6th to 9th rows show results on human objects and non-human objects.
@@
 OQ 51.3 31.3 17.4 51.6 31.1 17.3 47.5 26.5 26.0 48.1 21.2 30.7 48.1 18.3 29.6
-Appendix E Failure Cases
-We highlight two common failure scenarios encountered in our experiments.
-Firstly, although our model adeptly preserves contextual pixels, it struggles with accurately reconstructing non-target pixels within the mask area.
-This issue is illustrated in the first row of Figure 15, the position of ring in the source image is changed in the target image.
-The second challenge arises when dealing with subjects that exhibit a high degree of variability or freedom of movement.
-In such cases, as shown in the bottom row of the figure, accurately replicating the concept subject becomes difficult.
+Appendix 0.H Failure Cases
+Figure 14: Examples of failure cases. The model sometimes struggles to keep the details inside the mask area and could fail if the object has a high degree of freedom.
+We highlight one common failure scenario encountered in our experiments.
+The challenge arises when dealing with subjects that exhibit a high degree of variability or freedom of movement.
+In such cases, as shown in Fig.
+14, accurately replicating the concept subject becomes difficult.
 To address this, we are considering the implementation of explicit alignment, which we aim to explore in our future work.
-Figure 15: Examples of failure cases. The model sometimes struggles to keep the details inside the mask area and could fail if the object has a high degree of freedom.
-Appendix F Human Evaluation Interface
-Figure 16: The illustration of the user study interface.
-Amazon Turker was presented with one reference image mainly containing the concept subject, one source image to be swapped, and two generated images from SwapAnything and a baseline.